Project

General

Profile

NDOS Commissioning Meeting 2011/08/11

Attendance: Jon, Gary, Andrew, Denis, Leon

Agenda:

  1. Beam & Detector Update
  2. DataCheck Update
  3. DAQ Update
  4. DCS Update
  5. AOB

Beam & Detector Update

  • Enabled cooling on another 20 APDs. 1 of these (1-1-48) seems to be incapable of cooling. If any of these fail
  • Leon and Denis have put beads on TEC leads. Denis ran DSO scans on these (cooled), and remaining noise seems to have been killed an no new noise has been found. Gary: Nathan is interested in pursuing a software (firmware) fix for the noise; is it necessary? Leon: probably not, but worth pursuing.
  • Height of bosses where fibers are, and depth of groove ring, on optical connections are being measured on NDOS (on the bare FEBs). This seems to be affecting data taking. Leon: covers are being removed. Denis: power is not being disconnected from boxes.
  • Jon will call Chuck Andrews (CD facilities) to see if he is available to help us check for grounding issues. If he's not, perhaps we will do some basic tests of the system ourselves.
  • Peter: it doesn't look like the two views are in sync in the run that was just started. Experts are investigating.

DataCheck

  • No updated presentation, but:
  • Jan: moving php site to the nusoft area.
  • Fast processing is on hold since Jarek is at DPF.
  • Latest results from batch processing are in the development dB.

DCS Update

  • Athans:
    • IOC has correct scaling for all measured values
    • New version of GUI is in use; has alarmer capability
    • nova-daq-04 is being used by DCS experts
    • CSS is running on novadaq-master and novadcs-reader1. The one on novadaq-master will be moved to novadcs-master.
  • Martin:
    • found power/hv scripts in /usr/local/bin. These should not be there, Andrew N. will remove them.
    • procedure for rolling out new power/hv scripts? Jon: you decide, power really is a DCS issue.

DAQ Update

  • Peter: tried rolling out new message logger facility and again saw an issue with data logger. Immediately rolled back.

AOB

  • Jon: cancel night shifts? Consensus is yes. Night shifts will be cancelled through Nov. 13.